# RoSTE: An Efficient Quantization-Aware Supervised Fine-Tuning Approach for Large Language Models

### v2

@@ 5 Experiments
For the RoSTE algorithm, while we relaxed the lower level as a $\ell$-variable binary combinatorial problem (9), solving this sub-problem has a complexity of $\mathcal{O}(2^{\ell})$ which is still intractable for models like Llama 3.1 8B with $\ell=3\times 32+1$. As a remedy, we estimate the solution of (9) by computing only $\mathcal{E}({\cal W}^{kT},\{{\bf I}\}_{i=0}^{\ell-1})$ and $\mathcal{E}({\cal W}^{kT},\{{\bf H}\}_{i=0}^{\ell-1})$, then we determine each layer’s $\mathbf{R}_{i}$ by comparing the layer-wise quantization error. Lastly, we set $K=1$ where a one-shot rotation configuration adaptation by pre-trained model is found to perform well. We anticipate the performance to further improve with increasing $K$. More implementation details can be found in Appendices C, D.
Baselines. Besides the proposed RoSTE algorithm, we compare the performances of LLMs with quantized weight and activation obtained by two streams of baseline approaches. The first stream consists of applying PTQ methods on open-source supervised fine-tuned models in (Huang et al., 2024; Lambert et al., 2024). We reproduce the PTQ benchmarks using round-to-nearest (RTN) quantization, GPTQ (Frantar et al., 2022), knowledge distillation (LLM-QAT) (Liu et al., 2023), QuaRot (Ashkboos et al., 2024b) and SpinQuant (Liu et al., 2024). The second set consists of QAT methods, which apply STE and RoSTE to the SFT objective. The hyperparameters for reproducing our experiment results can be found in the Appendix at Table 4 and 5.
All experiments are conducted on a cluster of 8 NVIDIA A100 GPUs. Details of the training and evaluation settings can be found in Appendix C.
@@ 5.1 Accuracy of Quantized Models
For Exp.1, we present the accuracies of 4-bits (weights & activation) quantized, fine-tuned Pythia 6.9B found in Table 1. Herein, the best baseline is STE (without rotation). In comparison, RoSTE produces a quantized model that improves the average ROUGE score by +3.01. It recovers the performance of the full-precision SFT model with a gap of only -1.06 ROUGE score. For Exp.2, we present the accuracies of 4-bits (weights & activation) quantized, fine-tuned Llama 3.1 8B found in Table 2. Observe that RoSTE improved the average accuracy by +2.56 over the best baseline SpinQuant, despite there remains a gap of -10.47 below the full-precision fine-tuned model.
Lastly, in the appendix, we provide additional results on Pythia 1B and W4A8KV4 in Table 7, Qwen 2.5 in Table 8, and Llama 3.1 8B and W4A8KV4 quantization in Table 9.
@@ 6 Conclusion
This paper proposed the RoSTE algorithm for quantization-aware SFT training with an adaptive rotation strategy. To the best of our knowledge, this is one of the first algorithms for the efficient quantized SFT process. Besides achieving state-of-the-art performance, we also provide theoretical insights to justify the practical efficacy of RoSTE.
@@ Appendix F Additional Experiments
We show additional experiment results for the resultant accuracies of fine-tuning the LLMs with different configurations of quantization parameters. Particularly, the results for Exp.1 on the Pythia models can be found in Table 7, and the results for Exp.2 on the Pythia models can be found in Table 9. We observe consistent improvements with the RoSTE algorithm.
@@
Table 9: Additional experiments for Exp.2 on different bit-width configurations and different model sizes.

### v3

@@ 5 Experiments
For the RoSTE algorithm, while we relaxed the lower level as a $\ell$-variable binary combinatorial problem (9), solving this sub-problem has a complexity of $\mathcal{O}(2^{\ell})$ which is still intractable for models like Llama 3.1 8B with $\ell=3\times 32+1$. As a remedy, we estimate the solution of (9) by computing only $\mathcal{E}({\cal W}^{kT},\{{\bf I}\}_{i=0}^{\ell-1})$ and $\mathcal{E}({\cal W}^{kT},\{{\bf H}\}_{i=0}^{\ell-1})$, then we determine each layer’s $\mathbf{R}_{i}$ by comparing the quantization error layer-wise. Lastly, we set $K=1$ where a one-shot rotation configuration adaptation by pre-trained model is found to perform well. We anticipate the performance to further improve with larger $K$ on larger datasets. More implementation details can be found in Appendices C, D.
Baselines. Besides the proposed RoSTE algorithm, we compare the performances of LLMs with quantized weight and activation obtained by two streams of baseline approaches. The first stream consists of applying PTQ methods on open-source supervised fine-tuned models in (Huang et al., 2024; Lambert et al., 2024). We reproduce the PTQ benchmarks using round-to-nearest (RTN) quantization, GPTQ (Frantar et al., 2022) , QuaRot (Ashkboos et al., 2024b) and SpinQuant (Liu et al., 2024). The second set consists of QAT methods applied on the SFT objective, including STE and RoSTE. The hyperparameters for reproducing our experiment results can be found in the Appendix at Table 4 and 5.
All experiments are conducted on a cluster of 8 NVIDIA A100 GPUs. Details of the training and evaluation settings can be found in Appendix C. Statistics of the training cost (time, memory) can be found in Appendix G.
@@ 5.1 Accuracy of Quantized Models
For Exp.1, we present the accuracies of 4-bits (weights & activation) quantized, fine-tuned Pythia 6.9B and Qwen2.5 7B in Table 1. On quantizing Pythia 6.9B, the best baseline is STE (without rotation). In comparison, RoSTE produces a quantized model that improves the average ROUGE score by +3.01. It recovers the performance of the full-precision SFT model with a gap of only -1.06 ROUGE score. Similarly on Qwen2.5 7B model, RoSTE improves upon the best baseline by +1.78 ROUGE score, with a gap of -0.77 below the full-precision SFT model. For Exp.2, we present the accuracies of 4-bits (weights & activation) quantized, fine-tuned Llama 3.1 8B in Table 2. Observe that RoSTE improved the average accuracy by +2.56 over the best baseline SpinQuant, despite a gap of -10.47 below the full-precision fine-tuned model.
Lastly, in the appendix, we provide additional results of W4A4K4 and W4A8K4 quantization on Pythia 1B in Table 7, Qwen2.5 0.5B in Table 8, and Llama 3.1 8B in Table 9.
@@ 6 Conclusion
This paper proposed the RoSTE algorithm for quantization-aware SFT training with an adaptive rotation strategy. Besides achieving state-of-the-art performance, we also provide theoretical insights to justify the practical efficacy of RoSTE. To the best of our knowledge, this is the first algorithm that leverage adaptive rotation and the fine-tuning objective to produce an accurate quantized model.
Acknowledgements
M. Hong and Q. Wei are supported in part by NSF grants ECCS-2426064 and CIF-2414372. Q. Wei is also supported by an Amazon Machine Learning Systems Fellowship. C.-Y. Yau and H.-T. Wai are supported by the Shun Hing Institute of Advanced Engineering, CUHK, under Project #MMT-p5-23.
@@ Appendix F Additional Experiments
We show additional experiment results for the resultant accuracies of fine-tuning the LLMs with different configurations of quantization parameters. Particularly, the results for Exp.1 on the Pythia models can be found in Table 7, and the results for Exp.2 on the Pythia models can be found in Table 9. We observe consistent improvements with the RoSTE algorithm. Additional comparisons to QLoRA (Dettmers et al., 2023), LLM-QAT (Liu et al., 2023) and DuQuant (Lin et al., 2024) are provided in some setups, for instance, we observed a significant performance degradation on DuQuant when the KV cache is quantized below 8 bits in Table 8.
@@
W4A4KV8 QuaRot 31.96 10.98 24.73 24.88 23.13
DuQuant 33.47 12.13 25.28 25.30 24.05
Table 9: Additional experiments for Exp.2 on different bit-width configurations.
@@
Appendix G Statistics of Training Cost
Table 10 presents additional statistics for the training costs when using RoSTE and other benchmark algorithms. We observe that while achieving better performance, RoSTE requires only similar amount of computation costs compared to benchmarked algorithms.
Table 10: Training time and peak GPU memory consumption for obtaining a quantized fine-tuned Qwen2.5 7B from its pre-trained checkpoint on a server of 8 $\times$ A100.
Bit-width Method Training Time (hours) Peak Memory (GB)
FP16 SFT 2.1 300
LoRA $(r=64)$ 0.55 173
W4A4KV4 SFT $\rightarrow$ GPTQ 2.1 $\rightarrow$ 0 300 $\rightarrow$ 0
SFT $\rightarrow$ QuaRot 2.1 $\rightarrow$ 0 300 $\rightarrow$ 0
SFT $\rightarrow$ SpinQuant 2.1 $\rightarrow$ 1.3 300 $\rightarrow$ 263
QLoRA $(r=64)$ 0.83 98
STE 2.4 317
RoSTE 2.8 318